OpenStudy (anonymous):

A 9.35g sample of a solid was placed in a 15.00mL flask. The remaining volume was filled with Benzene in which the solid is insoluble. The solid and Benzene together weigh 17.33 g. The density of Benzene is 0.876 g/mL. What is the density of the solid? Please explain how to get to this answer.

OpenStudy (cuanchi):

1. calculate the weight of the benzene by subtracting the total weight minus the weight of the sample (17.33-9.35= 7.98 g) this is the mass of benzene 2. Then density= m/v, you can calculate the volume of benzene "v= mass/density" v= 7.98/0.876mL = 9.195mL 3. Subtract the volume of the benzene from the total volume to calculate the volume of the solid. 15.00-9.195= 5.805 mL 4. with the volume and the mass of the solid sample you can calculate the density of the solid.